OpenStudy (anonymous):

1+x^-1 over 1-x^-2 eliminate the complex fractions

OpenStudy (anonymous):

they are kind of endless, huh? \[\frac{1-x^{-1}}{1-x^{-2}}\] \[\frac{1-x^{-1}}{1-x^{-2}}\times \frac{x^2}{x^2}=\frac{x^2-x}{x^2-1}\]

OpenStudy (anonymous):

still not done though just like before

OpenStudy (anonymous):

\[\frac{x^2-x}{x^2-1}=\frac{x(x-1)}{(x+1)(x-1)}\] then cancel to finish

OpenStudy (anonymous):

Yea but is it the same because I put 1 + x^-1 and you put 1- x^-1

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

change the minus to a plus and get \[\frac{x^2+x}{x^2-1}=\frac{x(x+1)}{(x+1)(x-1)}=\frac{x}{x-1}\]

OpenStudy (anonymous):

Do you happened to know x^-1+y^-1 over x+y

hero (hero):

Begin with: \[\frac{\frac{1}{x} + \frac{1}{y}}{x + y}\] Re-write as: \[\left(\frac{1}{x} + \frac{1}{y}\right) \div{(x + y)}\] Reciprocate (x + y): \[\left(\frac{1}{x} + \frac{1}{y}\right) \times \frac{1}{(x + y)}\] Combine fractions 1/x and 1/y: \[\left(\frac{y}{xy} + \frac{x}{xy}\right) \times \frac{1}{(x + y)}\]\[\left(\frac{x+y}{xy} \right) \times \frac{1}{(x + y)}\] Multiply: \[\frac{(x+y)}{xy(x+y)}\] Cancel factors of 1: \[\frac{\cancel{(x+y)}}{xy\cancel{(x+y)}}\] Result: \[\frac{1}{xy}\]